OpenStudy (anonymous):

In a class of 30 students the mean age of 12 boys is 16years and 8 months and 16 girls mean age is 15 years 10months. Calculate the mean age in years and months.

OpenStudy (anonymous):

18* girls

OpenStudy (anonymous):

Number of months of total ages of boys = 12( 16x12+8)=2400 months Number of months of total ages of girls = 18( 15x12+10)=3420 months Total months of all students=2400+3420=5820 months What is the average age in moths?

OpenStudy (anonymous):

The average age in moths is 5820/30 =194 months How many years and month are in 194 months?

OpenStudy (anonymous):

194= 12 x 16+ 2 16 years and 2 months